# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ Abstract
Many-body localization (MBL) lends remarkable robustness to nonequilibrium phases of matter. Such phases can show topological and symmetry breaking order in their ground and excited states, but they may also belong to an anomalous localized topological phase (ALT phase). All eigenstates in an ALT phase are trivial, in that they can be deformed to product states, but the entire Hamiltonian cannot be deformed to a trivial localized model without going through a delocalization transition. Using a correspondence between MBL phases with short-ranged entanglement and locality preserving unitaries—called quantum cellular automata (QCA)—we reduce the classification of ALT phases to that of QCA. This method extends to periodically (Floquet) and quasiperiodically driven ALT phases, and captures anomalous Floquet phases within the same framework as static phases. The QCA framework further generalizes to include symmetry-enriched ALT phases (SALT phases), and provides a large class of soluble models suitable for realization in quantum simulators. In systematizing the study of ALT phases, we both greatly extend the classification of interacting nonequilibrium systems and clarify a confusion in the literature which implicitly equates nontrivial Hamiltonians with nontrivial ground states.
@@
Many-body localization (MBL) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] can protect robust, topological phases even in systems which are far from equilibrium [12, 13]. For instance, localized systems can support spontaneous symmetry breaking [12, 14, 15, 13], topological order [16, 17, 18, 19, 20, 21], and nontrivial phases with external driving [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], including time crystals [33, 34, 35, 36, 37, 38, 39, 40]. In fact, any Hamiltonian built out of commuting terms, of the kind frequently employed to study topological order [41, 42, 43, 44, 45, 46, 47, 48], becomes robustly localized by choosing strongly disordered coefficients for the commuting terms [20].
@@
Two MBL systems that can be deformed into one another without going through a delocalization transition belong to the same localized phase [24, 23, 49, 25, 29, 50, 31]. Two Hamiltonians certainly belong to distinct localized phases when they support distinct eigenstate order [12, 14, 15, 13]—their eigenstates cannot be deformed into one another without their correlation length diverging. However, in this Letter, we argue eigenstate order is not the right concept to obtain a general understanding of localized topological phases. For one thing, not all ground state topological phases are compatible with MBL. More importantly, there exist anomalous localized topological phases [^1] The label “anomalous” descends from Refs. [22, 24]. It is not related to other usages of the word “anomalous” in the context of topological phases of matter. (ALT phases) [22, 24, 23, 49, 31, 50]. ALT phases have no eigenstate order, but nonetheless cannot be deformed to a trivial model—in which the commuting terms have disjoint support [51]—without going through a delocalization transition.
@@
Figure 1: (a) Many-body localization is characterized by commuting LIOMs, $\tau^{z}_{j}$ [6]. The LIOMs of an MBL model with short-ranged entanglement can be prepared from trivial, single-site LIOMs $\sigma^{z}_{j}$ by a QCA, $\tau^{z}_{j}=V\sigma^{z}_{j}V^{\dagger}$. ALT phases cannot be deformed back to a trivial model without delocalizing, even though they have no eigenstate order. (b) Multi-tone driven ALT phases correspond to loops or tori of QCA. The $\Omega$-spectrum property $\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}$ (5) implies that $n$-tone-driven ALT phases in $d$ dimensions correspond to $(d-1)$-dimensional phases with $n-1$ drives. The translation QCA in $d=1$ [58] corresponds with the anomalous Floquet phase in $d=2$ [23], which in turn corresponds with a two-tone-driven phase in $d=3$ with circulating surface states.
@@
This Letter systematizes the study of ALT phases in strongly interacting systems. Specifically, we focus on localized topological phases satisfying a condition we refer to as short-ranged entanglement. We will argue that instead of eigenstate order, the right way to characterize all topological localized phases with short-ranged entanglement is via quantum cellular automata (QCA) [58, 59, 60, 56]—the locality preserving unitaries which map physical degrees of freedom to local integrals of motion (LIOMs) [Fig. 1(a)].
@@
As QCA in spin systems have been classified in one [58] and two [61] dimensions, this correspondence allows for a classification of ALT phases in low dimensional systems. The same formalism extends to periodically (Floquet [62]) and quasiperiodically driven ALT phases [31], which are classified by loops or tori of QCA, respectively. Thus, we can treat static and $n$-tone-driven ALT phases in $d$ dimensions on the same footing. We also considerably develop the study of the topology of QCA, which allows us to classify all ALT phases with $d-n\leq 2$ [Fig. 1(b)].
Symmetry-enriched ALT phases (SALT phases) can also be classified by QCA, and we compute the classification for any symmetry group $G$ and $d-n\leq 1$. Our methods naturally provide explicit soluble models of driven (S)ALT phases which, being finite depth circuits, are well suited to realization in quantum simulators [63, 64, 65].
@@
ALT phases.—We aim to classify many-body localized phases. A Hamiltonian $H$ is many-body localized when it possesses a complete set of commuting LIOMs $\tau^{z}_{j}$ such that [6]
@@
Here, $j$ indexes LIOMs, $\cdots$” indicates higher-body interactions, $h_{jk\ldots l}$ are real numbers which decay with the maximum distance between any of $\tau^{z}_{j},\ldots,\tau^{z}_{l}$, and we consider local degrees of freedom on the integer lattice in $d$ dimensions, ${\mathbb{Z}}^{d}$.
@@
MBL arising from strong random disorder is believed to be asymptotically unstable for $d>1$ [71], and doubt has arisen regarding the LIOM description even for $d=1$ [72, 73, 74, 75, 76, 77, 11]. The case of correlated disorder remains open [78, 79, 80, 81, 82]. Regardless, the Hamiltonian (1) continues to function as an effective model for astronomical prethermal timescales [72, 76, 77, 83, 10, 11]. Indeed, the timescale for topological response in a localized model tends to be parametrically separate from the thermalization time [31, 50, 23, 32, 25, 29, 30]. This justifies treating the topological classification of LIOM Hamiltonians (1) and the eventual thermalization process as separate problems. This Letter addresses the former.
@@
We say two MBL Hamiltonians $H_{0}$ and $H_{1}$ are in the same localized topological phase if there is a continuous path of MBL Hamiltonians $H_{t}$ ($t\in[0,1]$) which connects them. Ignoring long-range resonances in the spectrum of $H_{t}$, it is possible to construct a continuous path of locality preserving unitaries $V_{t}$ (that is, $V_{t}AV_{t}^{\dagger}$ is localized near the support of $A$ for all operators $A$) such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$ [84, 85, 86, 87, 88, 89]. Indeed, this unitary simply follows the adiabatic evolution of the many-body eigenstates. Including long-range resonances prevents an adiabatic construction of $V_{t}$ [90], but it is expected that sufficiently dilute resonances can be incorporated such that there still exists a $V_{1}$ that conjugates LIOMs of $H_{0}$ into those of $H_{1}$ and is continuously connected to the identity [9]. Thus, if $H_{0}$ and $H_{1}$ are in the same localized phase, their LIOMs are related by a continuous path of locality-preserving unitary operators. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@
An immediate consequence is that if $H_{0}$ and $H_{1}$ have distinct eigenstate order, they cannot be in the same localized phase. However, the converse is not true: even if they have the same eigenstate order, it need not be possible to deform $H_{0}$ to $H_{1}$. This phenomenon is exemplified by ALT phases, which have no eigenstate order, but which cannot be deformed to a trivial MBL model with LIOMs supported on a single lattice site. We denote these trivial LIOMs $\sigma^{z}_{j}$, even when the local Hilbert space dimension is not 2.
@@
An example of a nontrivial ALT phase is the three-dimensional 3-fermion Walker-Wang model [47, 48, 56]. In the absence of symmetry, this model is believed to have no eigenstate order [48, 91], but nonetheless it supports topologically ordered surface states protected by localization. Phrased in the language of MBL, [56, Theorem II.4] showed that the 3-fermion model, and, in fact, any MBL Hamiltonian $H$ with short-ranged entanglement—there are local unitaries $\tau^{x}_{j}$ which create and destroy excitations of $\tau^{z}_{j}$ without affecting other $\tau^{z}_{k}$ (Appendix A.1)—can be prepared from a trivial model by a QCA. That is, there is a QCA $V$ such that $H=VH_{\mathrm{trivial}}V^{\dagger}$ and $H_{\mathrm{trivial}}$ has single-site LIOMs $\sigma^{z}_{j}$ (Appendix A.2). We note that the property we call short-range entanglement for MBL Hamiltonians implies all eigenstates are short-range entangled in the usual sense.
Hence, short-range entangled MBL models can be parameterized by QCA as $VH_{\mathrm{trivial}}V^{\dagger}$. Paths of nondegenerate MBL Hamiltonians induce paths of unitaries as described above, so Hamiltonians in the same phase can be prepared by QCA which are connected by a path. Thus, the path-connected components of QCA can be used to classify phases of MBL Hamiltonians. An additional subtlety is that this parameterization is not unique—any QCA $V$ and $VP$ which differ by a permutation of the sites $P$ prepare identical LIOMs, and so represent Hamiltonians in the same phase (Appendix A.3).
@@
We further allow stabilization by the stacking (tensor product) of additional ancillae—$V_{0}$ and $V_{0}\otimes\mathbbm{1}$ are said to be stably equivalent, and are considered to prepare Hamiltonians in the same phase [58, 51, 92, 93]. The space of stable equivalence classes of $d$-dimensional QCA is denoted $\mathrm{QCA}_{d}$, and its connected components $\pi_{0}(\mathrm{QCA}_{d})$ form an abelian group under tensor products of QCA [58, 94]. Short-range entangled localized phases are classified by the quotient of this group by permutation QCA.
@@
The QCA formalism also lets us straightforwardly identify which phases are anomalous. For any simultaneous eigenstate $|0\rangle$ of all $\sigma^{z}_{j}$, $V|0\rangle$ is an eigenstate of $H$. This defines a map $e_{0}(V)=V|0\rangle$ from $\mathrm{QCA}_{d}$ to the space of invertible states [95] (short-range entangled states) in $d$-dimensions, $\mathrm{inv}_{d}$. There is an induced homomorphism
@@
Most examples of ALT phases in the literature are driven systems [24, 23, 49, 25, 29, 50, 31]. We specialize to the case of multi-tone time dependence [98, 28, 99] (which includes Floquet driving as a special case), such that the Hamiltonian can be parameterized $H(t)=H(\theta_{1}(t),\ldots,\theta_{n}(t))$, where $\theta_{k}(t)=\omega_{k}t$ is a drive phase defined modulo $2\pi$. It is convenient to assemble the drive phases into a vector ${\vec{\theta}}_{t}=(\theta_{1}(t),\ldots,\theta_{n}(t))$. Then we say that the system is MBL with short-ranged entanglement if it has a locality preserving generalized Floquet decomposition ($\hbar=1$[28, 99]

### v3

@@ Abstract
Many-body localization (MBL) lends remarkable robustness to nonequilibrium phases of matter. Such phases can show topological and symmetry breaking order in their ground and excited states, but they may also belong to an anomalous localized topological phase (ALT phase). All eigenstates in an ALT phase are trivial, in that they can be deformed to product states, but the entire Hamiltonian cannot be deformed to a trivial localized model without going through a delocalization transition. Using a correspondence between MBL phases with short-ranged entanglement and locality preserving unitaries—called quantum cellular automata (QCA)—we reduce the classification of ALT phases to that of QCA. This method extends to periodically (Floquet) and quasiperiodically driven ALT phases, and captures anomalous Floquet phases within the same framework as static phases. We considerably develop the study of the topology of QCA, allowing us to classify static and driven ALT phases in low dimensions. The QCA framework further generalizes to include symmetry-enriched ALT phases (SALT phases)—which we also classify in low dimensions—and provides a large class of soluble models suitable for realization in quantum simulators. In systematizing the study of ALT phases, we both greatly extend the classification of interacting nonequilibrium systems and clarify a confusion in the literature which implicitly equates nontrivial Hamiltonians with nontrivial ground states.
I Introduction
@@
Many-body localization (MBL) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] can protect robust, topological phases even in systems which are far from equilibrium [12, 13]. Localized systems can support spontaneous symmetry breaking [12, 14, 15, 13], topological order [16, 17, 18, 19, 20, 21], and nontrivial phases with external driving [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. In fact, any Hamiltonian built out of commuting terms, of the kind frequently employed to study topological order [41, 42, 43, 44, 45, 46, 47, 48], becomes robustly localized by choosing strongly disordered coefficients for the commuting terms [20].
Figure 1: The topological classification of gapped or MBL Hamiltonians aims to identify the path connected components of gapped or MBL models (respectively), viewed as a subspace of all local Hamiltonians. MBL Hamiltonians are a subspace of gapped Hamiltonians. The path component (in either classification) containing a specified trivial model is called the trivial phase. An anomalous localized topological phase (ALT phase) is a nontrivial MBL phase with no eigenstate order, and so in particular no ground state order, and thus is contained inside the trivial gapped phase. There are also nontrivial gapped phases with no MBL Hamiltonians.
Two MBL systems that can be deformed into one another without going through a delocalization transition belong to the same localized phase [24, 23, 49, 25, 29, 50, 31]. Two Hamiltonians certainly belong to distinct localized phases when they support distinct eigenstate order [12, 14, 15, 13]—their eigenstates cannot be deformed into one another without their correlation length diverging. Indeed, the lack of a delocalization transition throughout a deformation of the Hamiltonian implies all eigenstates retain only short-ranged correlations.
@@
However, in this Article, we argue eigenstate order is not the right concept to obtain a general understanding of localized topological phases. For one thing, not all ground state topological phases are compatible with MBL. More importantly, there exist anomalous localized topological phases [^1] The label “anomalous” descends from Refs. [22, 24]. Strictly, only the edges of these phases are anomalous in a similar sense to what is usually meant for topological phases of matter. (ALT phases) [22, 24, 23, 49, 31, 50]. ALT phases have no eigenstate order, but nonetheless cannot be deformed to a trivial model—in which the commuting terms have disjoint support [51]—without going through a delocalization transition.
The situation is illustrated in Fig. 1, with the example eigenstate being the ground state. Within the space of all local Hamiltonians, we identify those with a finite energy gap above the ground state. The topological classification of ground states aims to enumerate and describe all connected components (phases) of this space of gapped Hamiltonians. The phase containing a specified trivial model, consisting of a sum of single-site operators, is called the trivial phase. Among these gapped models, some are actually MBL. The topological classification of MBL phases, in complete analogy to gapped phases, aims to enumerate and describe all the connected components of this space of MBL Hamiltonians. However, there is not just one MBL phase within each gapped phase. There can be many, or none at all. Some or all of the nontrivial MBL phases which are contained within the trivial gapped phase are ALT phases.
@@
Figure 2: (a) Many-body localization is characterized by commuting LIOMs, $\tau^{z}_{j}$ [6]. The LIOMs of an MBL model with short-ranged entanglement can be prepared from single-site operators $\sigma^{z}_{j}$ by a QCA, $\tau^{z}_{j}=V\sigma^{z}_{j}V^{\dagger}$. ALT phases cannot be deformed back to a trivial model without delocalizing, even though they have no eigenstate order. (b) Multi-tone driven ALT phases correspond to loops or tori of QCA. The $\Omega$-spectrum property $\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}$ (9) implies that $n$-tone-driven ALT phases in $d$ dimensions correspond to (d1)(d-1)-dimensional phases with n1n-1 drives. The translation QCA in $d=1$ [58] corresponds with the anomalous Floquet phase in $d=2$ [23], which in turn corresponds with a two-tone-driven phase in $d=3$ with circulating surface states.
@@
This Article systematizes the study of ALT phases in strongly interacting systems. Specifically, we focus on localized topological phases satisfying a condition we refer to as short-ranged entanglement. We argue that instead of eigenstate order, the right way to characterize all topological localized phases with short-ranged entanglement is via quantum cellular automata (QCA) [58, 59, 60, 56]—the locality preserving unitaries which map physical degrees of freedom to local integrals of motion (LIOMs) [Fig. 2(a)] (Sec. II).
@@
As QCA in spin systems have been classified in one [58] and two [61] dimensions, this correspondence allows for a classification of ALT phases in low dimensional systems. The same formalism extends to periodically (Floquet [62]) and quasiperiodically driven ALT phases [31], which are classified by loops or tori of QCA, respectively. We argue that loops of QCA are completely classified through a picture of pumping lower-dimensional QCA to their boundary [Fig. 2(b)]. More technically, we conjecture that QCA form a mathematical structure known as an $\Omega$-spectrum [63, 64, 65]. This has wide-reaching implications for the classification of driven ALT phases, and in particular allows us to classify all $d$-dimensional ALT phases with $n$ incommensurate driving frequencies such that $d-n\leq 2$ (Sec. III).
Our methods extend immediately to symmetry-enriched ALT phases (SALT phases), which correspond to symmetric QCA [66, 26, 54, 67]. The $\Omega$-spectrum proposal provides a conjecture for the classification of SALT phases in terms of a generalized cohomology theory [63, 64, 65]. With this perspective, we compute the classification of $n$-fold loops of symmetric QCA (that is, the $n$th homotopy group) for any group $G$ and $d-n\leq 1$. This generalizes existing partial classifications for certain $G$ and $d=1$, $n=0$ [66, 26, 54, 67]. For these low dimensions, this calculation is not dependent on the $\Omega$-spectrum conjecture always being true (Sec. IV).
Finally, the pumping picture for QCA loops naturally provides explicit soluble models of driven ALT phases. These models reveal the anomalous edge dynamics of driven ALT phases, which cannot be realized in a short-range entangled MBL model of the same dimension as the edge. Similarly to short-ranged entangled phases of states, this anomalous edge dynamics is the generic signature of an ALT phase. Being finite depth circuits, these models are well suited to realization in current quantum simulators [68, 69, 70] (Sec. V).
II Anomalous localized topological phases
II.1 Localized phases
We aim to classify many-body localized phases. We use a strong notion of localization, namely the existence of a complete set of local integrals of motion (LIOMs). A localized phase is then defined as the set of LIOM Hamiltonians which can be deformed into one another while still possessing a complete set of LIOMs.
@@ II.1 Localized phases
More formally, a Hamiltonian $H$ is many-body localized when it possesses a complete set of commuting LIOMs $\tau^{z}_{j}$ such that [6]
@@
Here, $j$ indexes LIOMs, $\cdots$” indicates higher-body interactions, and $h_{jk\ldots l}$ are real numbers which decay with the maximum distance between any of $\tau^{z}_{j},\ldots,\tau^{z}_{l}$. For the sake of simplicity, we consider local degrees of freedom on the integer lattice in $d$ dimensions, ${\mathbb{Z}}^{d}$.
@@
MBL arising from strong random disorder is believed to be asymptotically unstable for $d>1$ [76], and doubt has arisen regarding the LIOM description even for $d=1$ [77, 78, 79, 80, 81, 82, 11]. The case of correlated (including quasiperiodic) disorder remains open [83, 84, 85, 86, 87]. Regardless, the Hamiltonian (1) continues to function as an effective model for astronomical prethermal timescales [77, 81, 82, 88, 10, 11]. Indeed, the timescale for topological response in the Hamiltonian (1) (which is finite in the LIOM model [31, 50, 23, 32, 25, 29, 30]) is parametrically separate from the thermalization time (which diverges exponentially as perturbations away from the LIOM model are reduced [11]). This vast parametric separation of time scales justifies treating the topological classification of LIOM Hamiltonians and the eventual thermalization process as separate problems. This Article addresses the former. We will henceforth assume the stability of MBL, with the understanding that in some cases this may be an approximate statement applying to an effective model.
We say two MBL Hamiltonians $H_{0}$ and $H_{1}$ are in the same localized topological phase if there is a continuous path of MBL Hamiltonians $H_{t}$ ($t\in[0,1]$) which connects them. We identify the trivial phase as the phase which contains an MBL Hamiltonian with LIOMs supported on single sites (Fig. 1). We denote these trivial LIOMs $\sigma^{z}_{j}$, even when the local Hilbert space dimension is not 2.
We seek to construct a continuous path of locality preserving unitaries $V_{t}$ such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$. A unitary $V$ is called locality preserving if $VAV^{\dagger}$ is localized near the support of $A$ for all local operators $A$. In a physical context, this means $VAV^{\dagger}$ has some operator decomposition where the norm of a term in the decomposition decays with its distance from the support of $A$. Any finite time evolution under a local Hamiltonian is locality preserving in this sense. Mathematically, it is convenient to deal with strictly locality preserving unitaries, where the support of $VAV^{\dagger}$ grows by a finite distance $r$. This includes all finite depth local quantum circuits. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@
Ignoring, for the moment, long-range resonances in the spectrum of $H_{t}$, the path of QCA $V_{t}$ we need may be constructed from the adiabatic evolution of the many-body eigenstates [89, 90, 91, 92, 93, 94]. Including long-range resonances prevents an adiabatic construction of $V_{t}$ [95], but it is expected that sufficiently dilute resonances can be incorporated such that there is still a $V_{1}$ that conjugates LIOMs of $H_{0}$ into those of $H_{1}$ and which is continuously connected to the identity. This is essentially the statement of the stability of MBL, as in Ref. [9]. Thus, if $H_{0}$ and $H_{1}$ are in the same localized phase, their LIOMs are related by a continuous path of QCA. Conversely, if the LIOMs of $H_{0}$ and $H_{1}$ are related by a path of QCA, it is straightforward to find a path between $H_{0}$ and $H_{1}$. Indeed, just set $H_{t}=V_{2t}H_{0}V_{2t}^{\dagger}$ for $t\in[0,1/2]$ and $H_{t}=2(1-t)V_{1}H_{0}V_{1}^{\dagger}+(2t-1)H_{1}$ for $t\in[1/2,1]$. The latter is localized as $V_{1}H_{0}V_{1}^{\dagger}$ and $H_{1}$ commute.
An MBL Hamiltonian is said to have eigenstate order if any of its eigenstates cannot be continuously deformed to a product state [12]. A familiar example is the toric code Hamiltonian with disordered coefficients [41]. This model is explicitly constructed from commuting terms—the LIOMs—and is well known to possess topological order.
The remainder of this article will restrict to MBL models with short-ranged entanglement. In the context of ground state phases, short-ranged entanglement means that all excitations on top of the state are topologically trivial. That is, there are local unitaries which can remove the excitations. We extend this notion to MBL models similarly. We say an MBL model has short-ranged entanglement if, for each LIOM $\tau^{z}_{j}$, there is a local unitary $\tau^{x}_{j}$ which can create and destroy excitations of $\tau^{z}_{j}$ without affecting other LIOMs $\tau^{z}_{k\neq j}$. We note that this property implies all eigenstates are short-range entangled in the usual sense. An excitation on top of any eigenstate can be removed by applying $\tau^{x}_{j}$ to all the excited LIOMs.
An example of a nontrivial short-ranged-entangled MBL model is the nontrivial fixed point of the Kitaev Majorana chain [96]. In contrast, while the toric code has LIOMs (the plaquette and star operators), the short-ranged entanglement condition is not met: the anyons can only be annihilated in pairs, and an isolated anyon cannot be removed locally [41]. Similarly, symmetry breaking phases are not short-ranged entangled, as domain walls cannot be annihilated locally.
II.2 Anomalous phases
Not all MBL phases are characterized by eigenstate order, as we describe in this section. Rather, the full classification of short-ranged entangled MBL phases can be obtained from the path-connected components of the space of quantum cellular automata in the same dimension, $\pi_{0}(\mathrm{QCA}_{d})$.
First, we note that eigenstate order, despite being a statement about individual states, is a robust property of the entire MBL Hamiltonian. This is an immediate consequence of the existence of the path of unitaries $V_{t}$ connecting LIOMs of Hamiltonians $H_{0}$ and $H_{1}$ in the same localized phase. Indeed, if $|\psi_{0}\rangle$ was an eigenstate of $H_{0}$, then $|\psi_{t}\rangle=V_{t}|\psi_{0}\rangle$ is a path of short-range correlated states demonstrating that $|\psi_{0}\rangle$ can be continuously deformed into an eigenstate $|\psi_{1}\rangle$ of $H_{1}$. Thus, $H_{0}$ and $H_{1}$ have the same eigenstate order. In contrapositive, Hamiltonians with distinct eigenstate order cannot be in the same localized phase.
@@ II.2 Anomalous phases
However, the converse is not true: even if they have the same eigenstate order, it need not be possible to deform $H_{0}$ to $H_{1}$. This phenomenon is exemplified by ALT phases, which have no eigenstate order, but which cannot be deformed to a trivial MBL model with LIOMs supported on a single lattice site.
@@
An example of a nontrivial ALT phase is the three-dimensional 3-fermion Walker-Wang model [47, 48, 56]. In the absence of symmetry, this model is believed to have no eigenstate order [48, 97], but nonetheless it supports topologically ordered surface states protected by localization. Phrased in the language of MBL, [56, Theorem II.4] showed that the 3-fermion model, and, in fact, any MBL Hamiltonian $H$ with short-ranged entanglement, can be prepared from a trivial model by a QCA. That is, there is a QCA $V$ such that $V^{\dagger}HV$ has single-site LIOMs $\sigma^{z}_{i}$. (We review the proof in Appendix A.) Note that, given an MBL Hamiltonian $H$, the choice of QCA $V$ is not unique. If we replace $V$ by $VP$, where $P$ is any QCA that maps $\sigma_{i}^{z}$ to a linear combination of products of $\sigma_{i}^{z}$’s (we refer to such QCA $P$ as generalized permutations, see Appendix A), then $V^{\dagger}HV$ also has $\{\sigma_{i}^{z}\}$ as LIOMs.
Moreover, using the property that paths of non-degenerate MBL Hamiltonians induce paths of unitaries that relate their LIOMs, as described above, it follows that if $H$ and $H^{\prime}$ are in the same short-range entangled MBL phase, then there exists a QCA $V^{\prime}$ which can be continuously connected to $V$ such that $(V^{\prime})^{\dagger}H^{\prime}V^{\prime}$ has $\{\sigma_{i}^{z}\}$ as LIOMs. We see that every short-range entangled MBL Hamiltonian can be prepared from $\{\sigma_{i}^{z}\}$ by a QCA, and those Hamiltonians in the same phase can be prepared by QCA in the same connected component. It follows that short-range entangled MBL phases can be classified by connected components of QCA (modulo generalized permutations).
@@
For the purposes of the classification, it is convenient to allow stabilization by the stacking (tensor product) of additional ancillae. The physical picture is that, for any system, there is an environment of additional degrees of freedom, and the classification of ALT phases (or MBL phases more generally) should be stable to the presence of such ancillae. Formally, the QCA $V_{0}$ and $V_{0}\otimes\mathbbm{1}$ are said to be stably equivalent, and are considered to prepare Hamiltonians in the same phase [58, 51, 98, 66]. The space of stable equivalence classes of $d$-dimensional QCA is denoted $\mathrm{QCA}_{d}$, and its connected components $\pi_{0}(\mathrm{QCA}_{d})$ form an abelian group under tensor products of QCA [58, 99]. Short-range entangled localized phases are classified by the quotient of this group by generalized permutation QCA.
@@
The QCA formalism also lets us straightforwardly identify which phases are anomalous. For any simultaneous eigenstate $|0\rangle$ of all $\sigma^{z}_{j}$, $V|0\rangle$ is an eigenstate of $H$. This defines a map $e_{0}(V)=V|0\rangle$ from $\mathrm{QCA}_{d}$ to the space of invertible states [100] (short-range entangled states) in $d$-dimensions, $\mathrm{inv}_{d}$. Our earlier observation that MBL Hamiltonians in the same phase have the same eigenstate order may be improved to the statement that there is an induced homomorphism from the classification of QCA to the classification of invertible states
@@
Ultimately, this homomorphism is just a mathematical formalization of the situation illustrated in Fig. 1. MBL phases are a subspace of gapped phases (and similarly, MBL phases with short-ranged entanglement are a subspace of gapped phases with short-ranged entanglement). The inclusion map from MBL phases to gapped phases induces a homomorphism between the classifications thereof, which is just to say that each MBL phase belongs to a definite gapped phase and the stacking operation on both sides of the map is the same. Some gapped phases have no MBL representatives, so that the homomorphism is not surjective. Further, distinct MBL phases might belong to the same gapped phase, so that the homomorphism is not injective.
II.3 Driven phases
Most examples of ALT phases in the literature are driven systems [24, 23, 49, 25, 29, 50, 31]. We specialize to the case of multi-tone time dependence [103, 28, 104] (which includes Floquet driving as a special case), such that the Hamiltonian can be parameterized as
$H(t)=H(\theta_{1}(t),\ldots,\theta_{n}(t)),$ (3)
where $\theta_{k}(t)=\omega_{k}t$ is a drive phase defined modulo $2\pi$. We show that short-range entangled MBL phases with $n$-tones can be classified using the higher homotopy groups of the space of QCA, denoted $\pi_{n}(\mathrm{QCA}_{d})$.
It is convenient to assemble the drive phases into a vector ${\vec{\theta}}_{t}=(\theta_{1}(t),\ldots,\theta_{n}(t))$. Then we say that the system is MBL with short-ranged entanglement if it has a locality preserving generalized Floquet decomposition ($\hbar=1$[28, 104]